Removing the Old Telescope

Daily events and images of the installation of the BBSO New Solar Telescope.

05 February Removing the cables from the old telescope is nearly completed. New vent louvers have been installed on the generator shed. The old generator is being replaced by a new instant-start generator. The Ramsey plate is being moved to over the door.
07 February The 10-inch and Singer telescope lenses were removed today.
08 February Work is also proceding on the dome, installing weather seals and working on the drive for the aperture cover. The aperture cover can now be opened, although the drive mechanism needs more work.
09 February Today we removed the 26 inch mirror, the top of the spectrograph, the Singer telescope and the electronics racks on the dome floor.
12 February Today we removed the 10 inch telescope.
14 February Roy is finding a new way to get into the dome.
20 February The 26-inch tank and the center counterweight assembly were removed from the dome today.
21 February The remiaining parts of the old telescope, the fork and polar axis, the pedestal, the 26-inch primary mirror, the Singer and 10-inch tubes and the spectrograph, were taken out of the dome today. The vacuum pump and hydraulic table were also removed.
22 February Old telescope parts to be disposed of, Didn't really want to take this picture.

© BBSO/NJIT - Last modified: February 22, 2007.